OpenStudy (anonymous):

Can someone help me out with this system? x+y=23 9x-7y=27

OpenStudy (abb0t):

I am sure you are familiar with method of substitution and or elimination, yES?

OpenStudy (anonymous):

yes @abb0t

OpenStudy (abb0t):

Which one are you more comfortable working with here?

OpenStudy (abb0t):

LOOOOL!

OpenStudy (anonymous):

@abb0t substitution is the one I am more comfortable with

OpenStudy (abb0t):

Ok. Well then here, this is what you will do you have two equations: \(\sf \color{red}{x+y=23}\) \(\sf \color{blue}{9x-7y=27}\) what you're going to do is rearrange one of those, i would suggest the one in red, to \(\sf \color{red}{y=23-x}\). notice that you have y = f(x). Take that f(x) - function and plug it into the blue equation! And solve for \(x\). \(\sf \color{blue}{9x-7(}\)\(\sf \color{red}{23-x}\)\(\sf \color{blue}{)=27}\) you know what to do now. distribute, solve for \(x\).

OpenStudy (anonymous):

@abb0t you are awesome! Thanks

OpenStudy (abb0t):

Once you solve for \(x\). You get a numerical value. That numerical valye that you get for \(x\). You can now use to solve for \(y\) in equation red.

OpenStudy (abb0t):

You can always check if your values are correct or not correct by plugging in the values for \(x\) and \(y\)! They should equal 23 or 27! If they do not, then you did something wrong. Good luck.